# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 6.C. End of the proof of Theorem D
Let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$ be a free product of groups, let =${\mathcal{F}}^{\prime}=\{[G_{1}],\ldots,[G_{k}]\}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$. We prove by induction on $k+N$ that $G\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
Suppose first that $k+N=1$. If $N=0$, then $G=G_{1}$ and $G\rtimes_{\Phi}\mathbb{Z}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by hypothesis. If $k=0$, then $G=\mathbb{Z}$, $G\rtimes_{\Phi}\mathbb{Z}$ is solvable and the result follows from [Weg15]. This proves the base case.
@@
By induction hypothesis, for every $[A]\in{\mathcal{F}}$, the group $A_{\Phi}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Combining the nonsporadic case (Lemma 6.1) and the sporadic case (Lemma 6.4), we conclude that $G\rtimes_{\Phi}\mathbb{Z}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. This concludes the proof. ∎
6.D. Proving A
We first record a corollary of D.
@@ 6.D. Proving A
Let $(G,{\mathcal{P}})$ be a virtually torsion-free relatively hyperbolic group with ${\mathcal{P}}$ finite and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. If for every $[P]\in{\mathcal{P}}$ we have $P_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
By 2.1 we may assume $G$ is torsion-free. Let ${\mathcal{F}}$ be the minimal free factor system of $G$ such that, for every $[P]\in{\mathcal{P}}$, there exists $[A]\in{\mathcal{F}}$ with $P\subseteq A$. Since $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$, by minimality of ${\mathcal{F}}$, we have $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$. Let $[A]\in{\mathcal{F}}$. We denote by ${\mathcal{P}}_{A}$ the peripheral structure of $A$ induced by ${\mathcal{P}}$. Since $G$ is torsion-free, the group $A$ is one-ended hyperbolic relative to ${\mathcal{P}}_{A}$. By Corollary 5.21 the group $A_{\Phi}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. By D, the group $G_{\Phi}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎
Finally, combining Corollary 5.21 and Corollary 6.5 proves our first theorem from the introduction.
@@
Let $(G,{\mathcal{P}})$ be a virtually torsion-free or one-ended relatively hyperbolic group with ${\mathcal{P}}$ finite and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. If for every $[P]\in{\mathcal{P}}$ we have $P_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Suppose $(G,{\mathcal{P}})$ is one-ended or virtually torsion free, and hyperbolic relative to finitely many conjugacy classes of slender subgroups. Then for every automorphism $\Phi$ of $G$, $\Gamma\coloneqq G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$.
@@
Knopf’s work on acylindrical actions of trees applies equally well in the setting of $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$ (see [Kno19, Corollary 4.2] and [BFW23, Theorem 2.4], note that Knopf does not state this but it is implicit in her work). We use the same trees as every step of the proof of A. Whenever a vertex group is identified as $\operatorname{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$, use 5.22 to further identify it as some $\operatorname{\mathrm{Fix}}(\Phi^{k})\rtimes_{\Phi}\mathbb{Z}$. As this has a finite index subgroup isomorphic to $\operatorname{\mathrm{Fix}}(\Phi^{k})\times\mathbb{Z}$, and $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$ passes to both subgroups and direct products, this vertex group lies in $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$. ∎
@@ 7. Proofs of the applications
Since for all $[P]\in{\mathcal{P}}$ the group $P$ is NRH we have that $\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$ is a finite index subgroup of $\operatorname{\mathrm{Aut}}(G)$. Let $\Phi\in\operatorname{\mathrm{Aut}}(G)$. The suspension $G_{\Phi}$ has a finite index subgroup $G_{\Phi^{n}}$ such that $\Phi^{n}\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. Now, A implies that $G_{\Phi^{n}}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. It follows from 2.1 that $G_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. The result now follows from 2.2. ∎
@@
Theorem 7.1 ([GL15, Theorem 4.3]).
Let $(G,{\mathcal{P}})$ be a relatively hyperbolic group. Suppose for every $[P]\in{\mathcal{P}}$, the group $P$ is finitely generated. If $G$ is one-ended relative to ${\mathcal{P}}$, then there is a short exact sequence
@@
By [BG06, Theorem 1.1] we see that $\operatorname{\mathrm{Out}}(G)$ is an arithmetic group. Hence, $\operatorname{\mathrm{Out}}(G)$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by [BFL14]. Technically they only prove the conjecture for $K$- and $L$-theory but it follows for $A$-theory by [Rüp16], [Kno19, Proof of Theorem 1.8(a)], and [${}^{+}$18, Theorem 6.19]. Alternatively, one may use [${}^{+}$18] and [KUWW18].
Since $G$ is virtually soluble it is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by [Weg15] (for $K$- and $L$-theory), [KUWW18] (for $A$-theory), and 2.1. Now, a virtually polycyclic group is poly-{virtually cyclic}, so any extension $G\rtimes\mathbb{Z}$ is also virtually polycyclic. Thus, $G\rtimes\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Further, note that $G/Z(G)$ is virtually polycyclic and so in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Combining these observations with 2.2 shows that $\operatorname{\mathrm{Aut}}(G)$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎
@@
This follows from the result for (orientable) surfaces [BB19] and assembling results in the literature. Note that Bartels–Bestvina only prove FJC for $K$- and $L$-theory but the result for $A$-theory follows (as usual) from [Kno19, Proof of Theorem 1.8(a)], and [${}^{+}$18, Theorem 6.19].
Let $S$ be a hyperbolic orbifold, and let $\Sigma$ be an orientable surface covering $S$ with finite degree so that $\pi_{1}(\Sigma)$ is characteristic in $\pi_{1}(S)$. (This can be achieved by taking any covering surface, passing to its orientation cover if necessary, and then taking the characteristic core of the corresponding subgroup and realising the covering surface.) By [KE21] there is an injective map from $\operatorname{\mathrm{Aut}}_{\text{geom}}(\pi_{1}(S))$ to $\operatorname{\mathrm{Aut}}_{\text{geom}}(\pi_{1}(\Sigma))$, where these geometric automorphism groups are exactly the lifts of the mapping class groups.
Restricting to the image, $\operatorname{\mathrm{Inn}}(\pi_{1}(S))$ will be normal, and by the third isomorphism theorem the quotient is isomorphic to $C/(\operatorname{\mathrm{Inn}}(\pi_{1}(S))/\operatorname{\mathrm{Inn}}(\pi_{1}(\Sigma)))$, where $C$ is a subgroup of $\operatorname{\mathrm{Aut}}_{\text{geom}}(\pi_{1}(\Sigma))/\operatorname{\mathrm{Inn}}(\pi_{1}(\Sigma))$, the mapping class group of $\Sigma$. The quotient of inner automorphism groups is finite (in fact isomorphic to the deck transformations $\pi_{1}(S)/\pi_{1}(\Sigma)$, so we have realised $\operatorname{\mathrm{MCG}}(S)$ as an extension
$1\to F\to C\to MCG(S)\to 1$
Since mapping class groups of surfaces are residually finite by [Gro74] (and residual finiteness passes to subgroups), we may apply Lemma 2.3 to obtain the conclusion. ∎
@@
If $G$ is a one ended group hyperbolic relative to finitely many conjugacy classes of virtually polycyclic groups, then $\operatorname{\mathrm{Out}}(G)$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
By 7.1 there is a finite index subgroup $\operatorname{\mathrm{Out}}_{0}(G)$ fitting into a short exact sequence
$1\to\mathfrak{T}\to\operatorname{\mathrm{Out}}_{0}(G)\to\prod_{i=1}^{p}\operatorname{\mathrm{MCG}}_{0}(S_{i})\times\prod_{j}\operatorname{\mathrm{Out}}(P_{j},\mathrm{Inc}_{P_{j}}^{(t)})\to 1$
We want to apply 2.2 to this short exact sequence. First we check the kernel $\mathfrak{T}$: this is a quotient of a direct product of virtually polycyclic groups, and hence is itself virtually polycyclic, and so in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by 7.2.
Now consider the image. The subgroups $\operatorname{\mathrm{Out}}(P_{j},\mathrm{Inc}_{P_{j}}^{(t)})$ are subgroups of $\operatorname{\mathrm{Out}}(P_{j})$ for a virtually polycyclic $P_{j}$, and hence are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by 7.2. Each $\operatorname{\mathrm{MCG}}(S_{i})$ maps with finite kernel onto the mapping class group of a hyperbolic 2-orbifold. By Lemma 2.3 it is enough to consider these mapping class groups. These are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by 7.3. Then the product is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by 2.1.
To apply 2.2 it remains to check the preimages of elements. These are of the form $\mathfrak{T}\rtimes\mathbb{Z}$, which are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by 7.2 since they are again virtually polycyclic. ∎
@@
Since ${\mathcal{P}}$ contains only finitely many conjugacy classes of virtually polycyclic groups, by [DS05, Corollary 1.14], we may modify ${\mathcal{P}}$ such that that for every $[P]\in{\mathcal{P}}$, the group $P$ is NRH. Observe that since $G$ is hyperbolic relative to finitely many conjugacy classes of virtually polycyclic subgroups its centre is finite. Hence, $\operatorname{\mathrm{Inn}}(G)$ is quasi-isometric to $G$ and again hyperbolic relative to finitely many conjugacy classes of NRH virtually polycyclic subgroups by [BDM09]. The result now follows from applying B to the short exact sequence
@@
C allows to prove that the outer automorphism groups of some small complexity relatively hyperbolic groups also belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
We prove the result for $\operatorname{\mathrm{Out}}(G)$, the proof for $\operatorname{\mathrm{Aut}}(G)$ being identical to the proof of C (this uses 8.3 when $C$ is nontrivial and A otherwise). Let $\operatorname{\mathrm{Out}}^{0}(G)$ be the index (at most) $2$ subgroup of $\operatorname{\mathrm{Out}}(G)$ preserving the conjugacy classes of $A$ and $B$. By [For02], every element $\phi\in\operatorname{\mathrm{Out}}^{0}(G)$ has a representative $\Phi\in\phi$ such that $\Phi(A)=A$ and $\Phi(B)=B$. Moreover, the map sending $\phi$ to $\Phi$ defines an isomorphism between $\operatorname{\mathrm{Out}}^{0}(G)$ and $\operatorname{\mathrm{Aut}}(A,C)\times\operatorname{\mathrm{Aut}}(B,C)$. By C, the groups $\operatorname{\mathrm{Aut}}(A,C)$ and $\operatorname{\mathrm{Aut}}(B,C)$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. By 2.1, the groups $\operatorname{\mathrm{Out}}^{0}(G)$ and $\operatorname{\mathrm{Out}}(G)$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎
@@
As above we only prove the result for $\operatorname{\mathrm{Out}}(G)$. Let $t$ be a stable letter for the HNN extension $A\ast_{C}$. By [Lev05], the group $\operatorname{\mathrm{Out}}(G)$ has an index $2$ subgroup $\operatorname{\mathrm{Out}}^{0}(G)$ such that any element $\phi\in\operatorname{\mathrm{Out}}^{0}(G)$ has a representative $\Phi\in\phi$ such that $\Phi(A)=A$ and $\Phi(t)=ta$ for some $a\in A$. Moreover, the map sending $\phi$ to $\Phi$ induces an isomorphism between $\operatorname{\mathrm{Out}}^{0}(G)$ and $A\rtimes\operatorname{\mathrm{Aut}}(A,C)$. Thus, $\operatorname{\mathrm{Out}}^{0}(G)$ fits in a short exact sequence
$1\to A\to\operatorname{\mathrm{Out}}^{0}(G)\to\operatorname{\mathrm{Aut}}(A,C)\to 1$
The group $\operatorname{\mathrm{Aut}}(A,C)$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by C. Moreover, for every infinite cyclic subgroup $Q\in\operatorname{\mathrm{Aut}}(A,C)$, the preimage of $Q$ in $\operatorname{\mathrm{Out}}^{0}(G)$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by A. Thus, by 2.2, the groups $\operatorname{\mathrm{Out}}^{0}(G)$ and $\operatorname{\mathrm{Out}}(G)$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎

### 02 Mar 2026 v2

@@ 6.C. End of the proof of Theorem D
Let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$ be a free product of groups, let ${\mathcal{F}}^{\prime}=\{[G_{1}],\ldots,[G_{k}]\}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$. We prove by induction on $k+N$ that $G\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
Suppose first that $k+N=1$. If $N=0$, then $G=G_{1}$ and $G\rtimes_{\Phi}\mathbb{Z}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by hypothesis. If $k=0$, then $G=\mathbb{Z}$, $G\rtimes_{\Phi}\mathbb{Z}$ is solvable and the result follows from [51]. This proves the base case.
@@
By induction hypothesis, for every $A]\in{\mathcal{F}}$, the group $A_{\Phi}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Combining the nonsporadic case (Lemma 6.1) and the sporadic case (Lemma 6.4), we conclude that $G\rtimes_{\Phi}\mathbb{Z}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. This concludes the proof. ∎
6.D. Proving ˜A
We first record a corollary of ˜D.
@@ 6.D. Proving ˜A
Let $(G,{\mathcal{P}})$ be a virtually torsion-free relatively hyperbolic group with $\mathcal{P}$} finite and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. If for every $P]\in{\mathcal{P}}$ we have $P_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
By ˜2.1 we may assume $G$ is torsion-free. Let $\mathcal{F}$} be the minimal free factor system of $G$ such that, for every $P]\in{\mathcal{P}}$, there exists $A]\in{\mathcal{F}}$ with $P\subseteq A$. Since $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$, by minimality of $\mathcal{F}$}, we have $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}})$. Let $A]\in{\mathcal{F}}$. We denote by ${\mathcal{P}}_{A}$ the peripheral structure of $A$ induced by $\mathcal{P}$}. Since $G$ is torsion-free, the group $A$ is one-ended hyperbolic relative to ${\mathcal{P}}_{A}$. By Corollary˜5.23 the group $A_{\Phi}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. By ˜D, the group $G_{\Phi}$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎
Finally, combining Corollary˜5.23 and Corollary˜6.5 proves our first theorem from the introduction.
@@
Let $(G,{\mathcal{P}})$ be a virtually torsion-free or one-ended relatively hyperbolic group with $\mathcal{P}$} finite and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. If for every $P]\in{\mathcal{P}}$ we have $P_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
Suppose $(G,{\mathcal{P}})$ is one-ended or virtually torsion free, and hyperbolic relative to finitely many conjugacy classes of virtually polycyclic subgroups. Then for every automorphism $\Phi$ of $G$, $\Gamma\coloneqq G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$.
@@
Knopf’s work on acylindrical actions of trees applies equally well in the setting of $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$ (see [37, Corollary 4.2] and [11, Theorem 2.4], note that Knopf does not state this but it is implicit in her work). We use the same trees as every step of the proof of ˜A. Whenever a vertex group is identified as $\operatorname{Per}(\Phi)\rtimes_{\Phi}\mathbb{Z}$, use ˜5.24 to further identify it as some $\operatorname{\mathrm{Fix}}(\Phi^{k})\rtimes_{\Phi}\mathbb{Z}$. As this has a finite index subgroup isomorphic to $\operatorname{\mathrm{Fix}}(\Phi^{k})\times\mathbb{Z}$, and $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$ passes to both subgroups and direct products, this vertex group lies in $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$. ∎
Remark $6.$
Note that in both ˜A and ˜E we may replace the hypothesis ‘virtually torsion free’ with the more general assumption that $G$ has a finite index subgroup whose Dunwoody decomposition has no non-trivial edge groups. We do not pursue this here.
@@ 7. Proofs of the applications
Since for all $P]\in{\mathcal{P}}$ the group $P$ is NRH we have that $\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$ is a finite index subgroup of $\operatorname{\mathrm{Aut}}(G)$. Let $\Phi\in\operatorname{\mathrm{Aut}}(G)$. The suspension $G_{\Phi}$ has a finite index subgroup $G_{\Phi^{n}}$ such that $\Phi^{n}\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. Now, ˜A implies that $G_{\Phi^{n}}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. It follows from ˜2.1 that $G_{\Phi}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. The result now follows from ˜$2.$ ∎
@@
Theorem 7.1 ([27, Theorem 4.3]).
Let $(G,{\mathcal{P}})$ be a relatively hyperbolic group. Suppose for every $P]\in{\mathcal{P}}$, the group $P$ is finitely generated. If $G$ is one-ended relative to $\mathcal{P}$}, then there is a short exact sequence
@@
By [9, Theorem 1.1] we see that $\operatorname{\mathrm{Out}}(G)$ is an arithmetic group. Hence, $\operatorname{\mathrm{Out}}(G)$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by [1]. Technically they only prove the conjecture for $K$- and $L$-theory but it follows for $A$-theory by [48], [37, Proof of Theorem 1.8(a)], and [18, Theorem 6.19]. Alternatively, one may use [18] and [36].
Since $G$ is virtually soluble it is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by [51] (for $K$- and $L$-theory), [36] (for $A$-theory), and ˜2.1. Now, a virtually polycyclic group is poly-{virtually cyclic}, so any extension $G\rtimes\mathbb{Z}$ is also virtually polycyclic. Thus, $G\rtimes\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Further, note that $G/Z(G)$ is virtually polycyclic and so in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Combining these observations with ˜2.2 shows that $\operatorname{\mathrm{Aut}}(G)$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎
@@
This follows from the result for (orientable) surfaces [2] and assembling results in the literature. Note that Bartels–Bestvina only prove FJC for $K$- and $L$-theory but the result for $A$-theory follows (as usual) from [37, Proof of Theorem 1.8(a)], and [18, Theorem 6.19].
Let $S$ be a hyperbolic orbifold, and let $\Sigma$ be an orientable surface covering $S$ with finite degree so that $\pi_{1}(\Sigma)$ is characteristic in $\pi_{1}(S)$. (This can be achieved by taking any covering surface, passing to its orientation cover if necessary, and then taking the characteristic core of the corresponding subgroup and realising the covering surface.) By [38] there is an injective map from $\operatorname{\mathrm{Aut}}_{\text{geom}}(\pi_{1}(S))$ to $\operatorname{\mathrm{Aut}}_{\text{geom}}(\pi_{1}(\Sigma))$, where these geometric automorphism groups are exactly the lifts of the mapping class groups.
Restricting to the image, $\operatorname{\mathrm{Inn}}(\pi_{1}(S))$ will be normal, and by the third isomorphism theorem the quotient is isomorphic to $C/(\operatorname{\mathrm{Inn}}(\pi_{1}(S))/\operatorname{\mathrm{Inn}}(\pi_{1}(\Sigma)))$, where $C$ is a subgroup of $\operatorname{\mathrm{Aut}}_{\text{geom}}(\pi_{1}(\Sigma))/\operatorname{\mathrm{Inn}}(\pi_{1}(\Sigma))$, the mapping class group of $\Sigma$. The quotient of inner automorphism groups is finite (in fact isomorphic to the deck transformations $\pi_{1}(S)/\pi_{1}(\Sigma)$), so we have realised $\operatorname{\mathrm{MCG}}(S)$ as an extension
$1\to F\to C\to MCG(S)\to 1.$
Since mapping class groups of surfaces are residually finite by [23] (and residual finiteness passes to subgroups), we may apply Lemma˜2.3 to obtain the conclusion. ∎
@@
If $G$ is a one-ended group hyperbolic relative to finitely many conjugacy classes of virtually polycyclic groups, then $\operatorname{\mathrm{Out}}(G)$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
By ˜7.1 there is a finite index subgroup $\operatorname{\mathrm{Out}}_{0}(G)$ fitting into a short exact sequence
$1\to\mathfrak{T}\to\operatorname{\mathrm{Out}}_{0}(G)\to\prod_{i=1}^{p}\operatorname{\mathrm{MCG}}_{0}(S_{i})\times\prod_{j}\operatorname{\mathrm{Out}}(P_{j},\mathrm{Inc}_{P_{j}}^{(t)})\to 1.$
We want to apply ˜2.2 to this short exact sequence. First we check the kernel $\mathfrak{T}$: this is a quotient of a direct product of virtually polycyclic groups, and hence is itself virtually polycyclic, and so in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by ˜7.2.
Now consider the image. The subgroups $\operatorname{\mathrm{Out}}(P_{j},\mathrm{Inc}_{P_{j}}^{(t)})$ are subgroups of $\operatorname{\mathrm{Out}}(P_{j})$ for a virtually polycyclic $P_{j}$, and hence are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by ˜7.2. Each $\operatorname{\mathrm{MCG}}(S_{i})$ maps with finite kernel onto the mapping class group of a hyperbolic 2-orbifold. By Lemma˜2.3 it is enough to consider these mapping class groups. These are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by ˜7.3. Then the product is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by ˜2.1.
To apply ˜2.2 it remains to check the preimages of elements. These are of the form $\mathfrak{T}\rtimes\mathbb{Z}$, which are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by ˜7.2 since they are again virtually polycyclic. ∎
@@
Since $\mathcal{P}$} contains only finitely many conjugacy classes of virtually polycyclic groups, by [17, Corollary 1.14], we may modify $\mathcal{P}$} such that that for every $P]\in{\mathcal{P}}$, the group $P$ is NRH. Observe that since $G$ is hyperbolic relative to finitely many conjugacy classes of virtually polycyclic subgroups its centre is finite. Hence, $\operatorname{\mathrm{Inn}}(G)$ is quasi-isometric to $G$ and again hyperbolic relative to finitely many conjugacy classes of NRH virtually polycyclic subgroups by [10]. By Proposition 7.4, the group $\operatorname{\mathrm{Out}}(G)$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. The result now follows from applying ˜B to the short exact sequence
@@
˜C allows us to prove that the outer automorphism groups of some small complexity relatively hyperbolic groups also belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
We prove the result for $\operatorname{\mathrm{Out}}(G)$, the proof for $\operatorname{\mathrm{Aut}}(G)$ being identical to the proof of ˜C (this uses ˜8.1 when $C$ is nontrivial and ˜A otherwise). Let $\operatorname{\mathrm{Out}}^{0}(G)$ be the index (at most) $2$ subgroup of $\operatorname{\mathrm{Out}}(G)$ preserving the conjugacy classes of $A$ and $B$. By [20], every element $\phi\in\operatorname{\mathrm{Out}}^{0}(G)$ has a representative $\Phi\in\phi$ such that $\Phi(A)=A$ and $\Phi(B)=B$. Moreover, the map sending $\phi$ to $\Phi$ defines an isomorphism between $\operatorname{\mathrm{Out}}^{0}(G)$ and $\operatorname{\mathrm{Aut}}(A,C)\times\operatorname{\mathrm{Aut}}(B,C)$. By ˜C, the groups $\operatorname{\mathrm{Aut}}(A,C)$ and $\operatorname{\mathrm{Aut}}(B,C)$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. By ˜2.1, the groups $\operatorname{\mathrm{Out}}^{0}(G)$ and $\operatorname{\mathrm{Out}}(G)$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎
@@
As above we only prove the result for $\operatorname{\mathrm{Out}}(G)$. Let $t$ be a stable letter for the HNN extension $A\ast_{C}$. By [39], the group $\operatorname{\mathrm{Out}}(G)$ has an index $2$ subgroup $\operatorname{\mathrm{Out}}^{0}(G)$ such that any element $\phi\in\operatorname{\mathrm{Out}}^{0}(G)$ has a representative $\Phi\in\phi$ such that $\Phi(A)=A$ and $\Phi(t)=ta$ for some $a\in A$. Moreover, the map sending $\phi$ to $\Phi$ induces an isomorphism between $\operatorname{\mathrm{Out}}^{0}(G)$ and $A\rtimes\operatorname{\mathrm{Aut}}(A,C)$. Thus, $\operatorname{\mathrm{Out}}^{0}(G)$ fits in a short exact sequence
$1\to A\to\operatorname{\mathrm{Out}}^{0}(G)\to\operatorname{\mathrm{Aut}}(A,C)\to 1.$
The group $\operatorname{\mathrm{Aut}}(A,C)$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by ˜C. Moreover, for every infinite cyclic subgroup $Q\in\operatorname{\mathrm{Aut}}(A,C)$, the preimage of $Q$ in $\operatorname{\mathrm{Out}}^{0}(G)$ belongs to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ by ˜A. Thus, by ˜2.2, the groups $\operatorname{\mathrm{Out}}^{0}(G)$ and $\operatorname{\mathrm{Out}}(G)$ belong to $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. ∎